Two Point Distortion and Rotation Theorems O. P. Juneja, Kanpur, India, S. Ponnusamy, Helsinki, Finland and S. Rajasekaran, Madras, India

Keywords: Univalent, bounded rotation, starlike and convex functions AMS Subject classification: 30C45, 30C55, 30A10 Abstract: This paper deals with two point characterization for functions in various sub families of univalent and non-univalent families which are defined in the unit disc.